### v3

# Blind Spot Navigation in Large Language Model Reasoning with Thought Space Explorer

Abstract
Recent advances in large language models (LLMs) have demonstrated their potential in handling complex reasoning tasks, which are usually achieved by constructing a thought chain to guide the model in solving the problem with multi-step thinking. However, existing methods often remain confined to previously explored solution spaces and thus overlook the critical blind spot within LLMs’ cognitive range. To address these issues, we introduce the “Thought Space Explorer” (TSE), a novel framework to expand and optimize thought structures to guide LLMs to explore their blind spots of thinking. By generating new reasoning steps and branches based on the original thought structure with various designed strategies, TSE broadens the thought exploration view and alleviates the impact of blind spots for LLM reasoning. Experimental results on multiple levels of reasoning tasks demonstrate the efficacy of TSE by surpassing various baseline methods. We also conduct extensive analysis to understand how structured and expansive thought can contribute to unleashing the potential of LLM reasoning capabilities.
Introduction
Recent advances in large language models (LLMs) have shown great potential in solving complex tasks with reasoning capabilities (Huang and Chang 2022; Patterson et al. 2022; Achiam et al. 2023; Mao et al. 2023; Dutta et al. 2025) by guiding the LLMs to reasoning logically solve the complex task step-by-step. A common practice is to design the Chain-of-Thought (CoT) (Kojima et al. 2022) to boost reasoning capabilities by evolving the thinking from single outputs to intermediate steps reasoning.
Existing studies (Wang et al. 2022; Yao et al. 2024; Zhang, Yuan, and Yao 2024; Besta et al. 2024; Pandita et al. 2025) attempt to develop various thought structures with multiple chains or branches of thought on top of CoT to arouse the LLM reasoning ability. Compared with direct output and CoT, the core advantage of thought structures enables models to explore the solution space of a task from local to global (Hao et al. 2023). For example, in Figure 1, thought structures may initiate exploration from two distinct points “specialty” and “industry”. Such exploration allows LLMs to generate diverse paths to solutions and thus enhances reasoning capacity. Moreover, the diverse structures can enable models to perform forward and backward evaluations within the explored thought space to seek the optimal solution, i.e., a more effective reasoning thought.
A series of studies are conducted to optimize thought structures with various aspects, including generating parallel thought (Wang et al. 2022), constructing tree structure reasoning topology on top of CoT (Yao et al. 2024), fine-tuning the LLMs with direct preference optimization (DPO) to align thought steps of CoT with task performance (Zhang et al. 2024c), etc. The key idea of this work (see Figure 1) is to compare multiple responses or extending existing chains (e.g. “Coffee over Drone” or “Coffee industry $\rightarrow$ Coffee bottle industry”) to obtain a better thought chain.
However, these approaches lack exploration of unknown solution spaces, which might lead to consistent oversight of LLMs’ cognitive blind spots (Zhang et al. 2024a; Sprague et al. 2025; Mo et al. 2024; Liu et al. 2025; Weerasooriya et al. 2023). Merely generating more chains does not enable LLMs to conceive of content previously unthought of. As described in Figure 1, over-generated chains tend to repeat prior thought patterns, leading to two main issues: (1) the absence of feasible solutions—when such solutions lie in blind spot regions, repeatedly filtering or extending existing paths may converge to a local optimum (e.g., exploring only from a coffee perspective); and (2) insufficient diversity—especially for open-ended questions, where existing methods have limited impact on exploring the thought space, and excessive extension or filtering might even reduce the diversity of responses (e.g., discarding feasible solutions or creating redundancy through repetitive thinking).
To address these issues, we propose the Thought Space Explorer (TSE), a novel framework designed to expand and optimize thought structures. The TSE starts from thought paths already explored and guides the model to explore hidden solution spaces because the existing thought structures often already contain feasible solutions or crucial information pointing towards such solutions. To enhance efficiency and precision, further exploration of the model starts from thought nodes within explored solutions, which ensures that the reasoning process is not a blind exploration but a deeper inquiry based on verified insights.
Figure 1: Thought structure optimization through TSE. On the left side, we showcase traditional thought structures and optimization methods, where the LLMs’ generation may limited by its thought pattern. On the right side, we show how TSE expands thought structure through a three-step generation of branches. TSE guides LLMs to explore the blind spots between previous thought paths.
To identify key points of information from existing thoughts, as shown in Figure 1, we first quantify each thought node’s contribution to the conclusion during the model’s reasoning process to select key nodes (e.g., in Chain 2, the details about “drones and delivery” serve as key information leading toward “logistics industry”). Considering the visibility of parameters in LLMs, we select key nodes from two perspectives: relative gradients and semantic relationships. Based on these key nodes, the model then generates new thought nodes and proceeds with deeper reasoning in new directions from “original nodes” to “new nodes”, facilitating exploration of the solution space through the thought structure. Finally, depending on whether the reasoning task requires a singular or comprehensive conclusion, we perform collaborative reasoning across the entire thought structure to generate the output.
We evaluate the effectiveness of TSE on three reasoning tasks and the results show that TSE significantly improves the performance of thought structures compared with existing methods. We further analyze the effectiveness of each component to comprehensively understand their functionality in each stage.
Our contributions include: (1) We propose TSE reasoning framework to expand thought structures for exploring solution spaces to alleviate the impact of blind spots for LLM reasoning. (2) We investigate various strategies to prioritize and refine the thought structure by identifying the importance of the node in the thought structure. Different strategies are designed to adapt to various settings. (3) Experimental results on three specific reasoning tasks indicate the effectiveness of our TSE compared with the existing reasoning methods without exploring thought structure.
Related Work
LLM Reasoning Structures
The most straightforward method to address reasoning tasks is to generate a conclusion through one-step thinking. However, the LLM might overlook essential intermediate steps and generate incoherent logic or incorrect conclusions (Chu et al. 2023). The advent of CoT (Wei et al. 2022; Wang and Zhou 2024) optimizes the reasoning step by connecting distinct thoughts into a coherent sequence (Li et al. 2024). Although CoT can improve transparency and coherence, its singular structure limits its capability to handle more complex logical relationships (Jin and Lu 2024; Sprague et al. 2025). To this end, some studies develop structured reasoning methods, such as self-consistent CoT and Tree-of Thought. (Wang et al. 2022; Zhang, Yuan, and Yao 2024; Yao et al. 2024; Liu et al. 2024; Mo and Xin 2024; Zhang et al. 2024b, a). These sophisticated thought structures enhance the consistency and systematic nature of reasoning by expanding the ability of the model to manage diverse logical relationships (Xia et al. 2024; Stechly, Valmeekam, and Kambhampati 2024). Thought structures offer distinct advantages by maintaining coherence and depth while increasing the diversity and flexibility of reasoning paths (Liang et al. 2024). However, the reasoning chains within these structures are highly repetitive and thus reduce generation and selection efficiency.
Thought Structure Optimizations
To further enhance the capabilities of thought structures, recent researches focus on two main optimization strategies. The first is selecting optimal paths within the structures (Feng et al. 2023; Long 2023; Hao et al. 2023; Shinn, Labash, and Gopinath 2023; Jung et al. 2022). By choosing optimal paths, the model filters the irrelevant and low-quality branches and globally searches for correct or optimal solutions, thus enhancing reasoning efficiency and accuracy. The second is expanding reasoning depth and breadth (Zhu et al. 2022; Besta et al. 2024; Gao et al. 2024; Zhang et al. 2024b, c; Hou et al. 2024; Zhang et al. 2025). By deepening and widening the thought structure, the model can explore a broader array of possibilities and perspectives, thus improving its understanding and capability to solve complex issues (Chen et al. 2025; Pan et al. 2025; Wang et al. 2025). However, these methods might be limited to previously explored spaces and directions, thus failing to adequately investigate the blind spots within the thought space of the model. Different from them, we aim to expand the depth and breadth of thought structures by actively exploring the cognitive blind spots of the model in thought space.
Methodology
To expand and optimize thought structures for effective exploration of reasoning spaces, we introduce TSE, a self-expansion and exploration method that allows language models to proactively address deficiencies in reasoning processes and explore new reasoning directions with limited steps of generation. In a step-by-step reasoning scenario, we implement the TSE method through several steps: (1) Key Node Selection and Generation, in which we select the nodes that are most influential to exploration directions based on the crucial information contained previously, then a new node is generated to integrate the insights of the key node for new exploration directions; (2) Connection and Expansion, in which we connect and expand the exploration to potential new directions within the semantic context from the newly-generated node; and (3) Collaborative Reasoning, in which we address deficiencies in the model’s ability to synthesize and integrate diverse reasoning paths in different directions.
Problem Formulation
Given a specific reasoning task $\mathcal{Q}$, we apply a large language model (LLM) $\mathcal{L}$ to a structured reasoning process as a thought structure $\mathcal{S}$. This structure consists of multiple reasoning sentences as thought nodes that are connected sequentially. We define the set of all thought nodes as $\mathcal{T}$, where each node $\mathcal{T}_{ij}$ represents the $j$-th reasoning step in the $i$-th thought chain. The thought structure $\mathcal{S}$ can be viewed as a directed graph consisting of vertices (thought nodes) $\mathbf{V}$ and edges (connections between consecutive nodes) $\mathbf{E}$, and a thought chain $C_{i}$ as an ordered sequence of thought nodes as
$\mathbf{V}$ $=\bigcup_{i=1}^{N}\bigcup_{j=1}^{K_{i}}\{T_{ij}\},\quad K_{i}=|C_{i}|,$ (1)
$\mathbf{E}$ $=\bigcup_{i=1}^{N}\bigcup_{j=1}^{K_{i}-1}\{(T_{ij},T_{i(j+1)})\}$
where $N$ and $K$ is the number of the thought chain $C$ and the number of the node $\mathbf{V}$ contained in a thought chain $C$, respectively.
$\mathcal{S}=(\mathbf{V},\mathbf{E}),\quad C_{i}=\langle T_{i1},T_{i2},\ldots,T_{iK_{i}}\rangle$ (2)
Figure 2: Illustration of solution space exploration via our TSE method. By generating new branches of solutions, the explored space of solutions expands.
For a specific task $\mathcal{Q}$, the complete reasoning solution space $\mathcal{P}$ encompasses all possible valid reasoning paths $C$ (thought chains). As shown in Figure 2, the space that has been explored by the generated thought structure $\mathcal{S}$ and the unexplored implicit thought space are denoted as $\mathcal{P_{S}}$ and $\mathcal{P_{U}}$, where $\mathcal{P_{S}}\cup\mathcal{P_{U}}=\mathcal{P}$.
Our goal is to actively expand the thought structure $\mathcal{S}$ by generating new reasoning branches $C^{\prime}$ to explore previously untouched areas of the solution space ($\mathcal{P_{U}}$). In this way, we maximize the chance of discovering effective and innovative solutions. The optimization goal is defined as $\max_{\mathcal{S}^{\prime}}J(\mathcal{S}^{\prime},\mathcal{Q})$, where $J$ is the reasoning performance metric and $\mathcal{S}^{\prime}$ is the expanded thought structure.
Key Node Selection and Generation
First, we select the most impactful nodes from the existing thought structure $\mathcal{S}$ for expansion. Specifically, we expect the nodes to contain crucial information necessary for solving the given problem by achieving the following aspects: (i) effective exploration of promising regions of the solution space by starting from these key nodes, thereby increasing the likelihood of discovering viable solutions; and (ii) minimization of error propagation by conducting additional analysis and verification on these critical nodes, which often represent potential sources of mistakes. To support the different availability of the backbone models, we propose two methods, gradient-based and self-prompting selection to identify the key nodes for those with accessible gradients and black-box ones, respectively.
Gradient-based Selection
The gradient-based selection is applicable when the internal structure and the gradient information of the model $\mathcal{L}$ are accessible. The representation of each thought node $\mathcal{T}_{ij}$ in the generated thought chains is obtained via the model $\mathcal{L}$ as:
$\mathcal{L}:\mathcal{T}_{ij}\rightarrow\mathbf{v}_{ij}\in\mathbb{R}^{d}$ (3)
Then, we aim to analyze the gradient importance of nodes $\mathcal{T}_{ij}$ relative to the conclusion node $\mathcal{T}_{iK_{i}}$ within a chain. The representation of the conclusion node $\mathbf{v}_{iK_{i}}$ is mapped to the output space as the model’s prediction at the end of the chain:
$\hat{y}_{i}=f(\mathbf{v}_{iK_{i}})$ (4)
where the function $f(\cdot)$ represents a mapping from the representation space of the conclusion node $\mathbf{v}_{iK_{i}}$ to the output space, and the prediction $\hat{y}_{i}$ is usually a textual conclusion or decision in terms of task $\mathcal{Q}$.
The self-information loss $L_{i}$ is a common practice to evaluate the model’s confidence in its predictions (Wang and Feng 2021), where higher confidence corresponds to lower loss values. Thus, we calculate the partial derivative of the loss $\mathbf{g}_{ij}$ with respect to each node’s representation $\mathbf{v}_{ij}$ and the Euclidean norm of its gradient $G_{ij}$ to measure the importance of the nodes. Then, we apply a normalization to determine the relative importance $I_{ij}$ of each node for a consistent and comparative analysis of node importance across different chains within the thought structure $\mathcal{S}$:
$L_{i}=-\log P(\hat{y}_{i}\mid\mathbf{v}_{iK_{i}}),\quad\mathbf{g}_{ij}=\frac{\partial L_{i}}{\partial\mathbf{v}_{ij}}$ (5)
$\quad G_{ij}=\|\mathbf{g}_{ij}\|_{2},\quad I_{ij}=\frac{G_{ij}}{\sum_{k=1}^{K_{i}}G_{ik}}$ (6)
In this way, we can identify the key nodes that have the highest impact on the model’s predictions. Now we have obtained a set of key nodes $\mathcal{T}_{key}=\left\{T_{i_{key}}\right\}_{i=1}^{N}$ with the highest relative gradient.
Self-prompting Selection
When the gradient is not accessible, e.g., a black-box model, the self-prompt selection is an alternative for key node identification. Under this setting, we leverage LLM’s natual language understanding capabilities to analyze and prioritize key nodes in the thought chains. Although the inner mechanisms of the model are opaque, we infer critical areas within the network’s structure by constructing specific prompts based on semantic and logical relationships. Specifically, model $\mathcal{L}$ ranks the importance of nodes within chain $C_{i}$ for the key node as:
$\mathcal{T}_{key}=\arg\max_{T_{ij}\in C_{i}}\text{rank }\mathcal{L}(T_{ij}\mid\mathcal{Q})$ (7)
Generating New Nodes
With the selected key nodes, the next step is to adopt them as conditional information to generate new thought nodes. Then we generate the new nodes based on the key nodes by integrating them into the thought structure and expanding them to form new branches. For each new node, we select two key nodes from $\mathcal{T}_{key}$, denoted as $\mathcal{T}_{i_{key}}$ and $\mathcal{T}_{l_{key}}$. With each pair of selected key nodes, the model generates a new thought node $\mathcal{T}^{1}_{il}$ as:
$\mathcal{T}^{1}_{il}=\mathcal{L}(\mathcal{T}_{i_{key}},\mathcal{T}_{l_{key}}),i,l\in[1,N],i\neq l$ (8)
Connection and Expansion
In terms of a new node, to decide which key node it should follow, we need to select $\mathcal{T}_{i_{key}}$ or $\mathcal{T}_{l_{key}}$ to serve as the connection point $\mathcal{T}^{1}_{il}$ for extending the new thought branch, which is supposed to maintain logical coherence while exploring previously unknown regions of the thought space. Therefore, we select the conditional nodes with higher semantic relevance with new nodes and contribute more to reasoning as connection nodes. Building on the previous stage, we propose two connection node selection methods, including the relative gradient method and the semantic relevance method.
Relative Gradient Selection
The relative gradient selection is applied to take the node between $\mathcal{T}_{i_{key}}$ and $\mathcal{T}_{l_{key}}$ with the larger relative gradient as the connection point by comparing the importance indices $I_{i_{key}}$ and $I_{l_{key}}$ as
$C^{\prime}=\langle\arg\max_{\mathcal{T}_{key}\in\{\mathcal{T}_{i_{key}},\mathcal{T}_{l_{key}}\}}(I_{key}),\mathcal{T}^{1}_{il}\rangle$ (9)
The $C^{\prime}$ denotes the new branch initiated from the key node with the higher importance index. Starting from $\mathcal{T}^{1}_{il}$, the model $\mathcal{L}$ continues conduct the reasoning guided by the information from the new nodes, and since the new nodes introduce fresh perspectives of existing information, the new branch is likely to go and explore previously unconsidered directions with step-by-step reasoning. The process of extending the branch $C^{\prime}$ continues until a specified depth is reached. Eventually, the language model $\mathcal{L}$ recursively generates subsequent nodes, using the strongest connection node as the context for generating new nodes as
$C^{\prime}$ $\rightarrow\mathcal{L}(C^{\prime})=\langle C^{\prime},\mathcal{T}^{2}_{il},\ldots,\mathcal{T}^{K}_{il}\rangle,\quad$ (10)
[…]
Semantic Relevance Selection
Semantic relevance selection is applicable after both key node selection methods. The connection node is selected by the model $\mathcal{L}$ based on semantic relationships between nodes. Then the new branch $C^{\prime}$ continues to extend recursively until a specified depth is reached, with $\mathcal{L}$ generating subsequent nodes based on the strongest semantic relationship as:
$C^{\prime}\rightarrow\mathcal{L}(C^{\prime})=\left\langle T_{\text{selected}},\mathcal{T}_{il}^{1},\mathcal{T}_{il}^{2},\ldots,\mathcal{T}_{il}^{K}\right\rangle$ (11)
where $\mathcal{T}_{\text{selected}}$ is the key node chosen by $\mathcal{L}$, and $K$ denotes the ending depth of the thought chain.
Collaborative Reasoning
Give the task $\mathcal{Q}$ and its unseen complete solution space $\mathcal{P}$, the model $\mathcal{L}$ needs to generate new thought branches based on the original thought structure $\mathcal{S}$. During this process, each new branch $C^{\prime}$ expands the explored thought space $\mathcal{P_{S}}\in\mathcal{P}$ by mining potential solutions based on the established structure:
$\mathcal{P_{S}^{\prime}}\leftarrow\mathcal{P_{S}}\cup\{C^{\prime}\},\quad\mathcal{P_{U}^{\prime}}\leftarrow\mathcal{P}-\mathcal{P_{S}^{\prime}}$ (12)
Thus the refined structure $\mathcal{S^{\prime}}$ with more solutions compared with $\mathcal{S}$ achieves a larger explored thought space $|\mathcal{P_{S}^{\prime}}|\geq|\mathcal{P_{S}}|$. Based on $\mathcal{S^{\prime}}$, we collaborate between original and new thought paths to obtain a unified reasoning conclusion.
Collaborative Weighted Summation
This strategy applies to gradient-available LLMs. We encompass all thought chains in the refined thought structure $\mathcal{S}^{\prime}$. First, we use gradient information to recalculate and select key nodes of each chain. For each key node $\mathcal{T}_{i_{key}}$, we calculate a weight $w_{i_{key}}$ based on its impact on the solution as:
$w_{i_{key}}=\frac{\exp(-L_{i_{key}})}{\sum_{k=1}^{K^{\prime}_{i}}\exp(-L_{ik})}$ (13)
where $L_{i_{key}}$ represents the self-information loss at node $\mathcal{T}_{i_{key}}$, which reflects the model’s confidence and potential errors in its prediction at that node. The contribution of each key node, denoted as $\mathbf{v_{i_{key}}}$, is the value vector that quantifies the influence of the node on the overall reasoning process. This vector can include various factors such as the relevance of the node’s content to the task or the accuracy of its inference. Then, for the reasoning task $\mathcal{Q}$, we compute the collaborative reasoning score by summing the contributions of all key nodes across all chains as:
$C(\mathcal{Q})=\sum_{i=1}^{N}\sum_{j=1}^{K_{i}}w_{i_{key}}\cdot\mathbf{v}_{ij}$ (14)
This process integrates the weighted contributions for a reasoning score that involves the individual node’s direct impact and its significance within the context of the entire chain. The final decision $D$ for the task $\mathcal{Q}$ is selected based on the highest collaborative reasoning score via
$D=\arg\max_{q\in\mathcal{Q}}C(q)$ (15)
LM-as-a-Judge
This strategy is suitable for reasoning tasks that require detailed interpretation and judgment, applicable to both gradient-available models and black-box models. Here, the model $\mathcal{L}$ acts as a judge to score each thought chain’s output based on its assessment of reasoning coherence, prediction confidence, or relevance to the task. Based on the evaluation, the model employs a voting mechanism to determine the final output:
$\text{Score}_{i}=\text{Evaluate}_{\mathcal{L}}(C_{q}),\quad D=\arg\max_{q\in\mathcal{Q}}C(q)$ (16)
In this way, we implement distinct collaboration strategies that enhance the model’s reasoning capabilities by leveraging targeted evaluations and adaptive integration.
Experiments
Task Game of 24-GPT Creative Writing-GPT
Success Rate (%) Soundness Innovation Coherence Expression Overall
CoT 13.30 5.26 5.24 5.03 4.95 5.12
CoT-SC 46.70 5.41 5.17 5.17 5.16 5.23
ToT 52.70 5.90 5.36 5.40 5.43 5.52
RATT 41.30 6.02 6.22 5.69 5.67 5.90
TSE 74.00 6.19 6.55 5.74 5.64 6.03
Task Game of 24-Llama Creative Writing-Llama
Success Rate (%) Soundness Innovation Coherence Expression Overall
CoT 1.33 4.63 3.98 4.25 4.20 4.27
CoT-SC 5.33 4.84 4.35 4.51 4.46 4.54
ToT 8.00 4.86 4.95 4.78 4.65 4.81
RATT 8.67 5.10 4.97 4.89 4.97 4.98
TSE 18.75 5.03 5.11 5.05 5.15 5.09
Table 1: Overall performance on reasoning tasks based on GPT-4o-mini and Llama. Bold and underline indicate the best and the second-best results.
Experimental Settings
Implementation Details and Baselines. We conduct experiments on three reasoning tasks that require mathematical, hierarchical, and comprehensive reasoning abilities to evaluate the effectiveness of our TSE method. We conduct these experiments utilizing the GPT-4o-mini (OpenAI 2024) and Llama-3.1-8B-Instruct (Hugging Face 2024). Unless otherwise specified, we generate five parallel thought chains with a depth of 5 for each question and use this structure as the basis for new node generation. The temperature for all models is set to the default value of 0.7, with a maximum token limit of 50. All tasks are performed on an NVIDIA 4090 GPU. We apply our method to the simplest multi-chain thought structure and compare the results with several baseline methods, including CoT, Vanilla CoT-SC, ToT, and RATT. This comparison aims to illustrate how our method enhances the thought structure compared to existing approaches.
Task Description. We evaluate TSE and the baseline methods on three task-specific reasoning datasets: (1) Game of 24 (Yao et al. 2023), a mathematical challenge to use the four basic arithmetic operations to make four given numbers equal 24. The task requires the language models to combine multiple operations to achieve a target outcome, which can evaluate arithmetic reasoning and logical problem-solving capabilities. (2) Mini Crosswords (Yao et al. 2023), a game of 5×5 mini crosswords and each input includes the 5 horizontal and 5 vertical clues. To solve this task, the model requires deeper exploration and strategic integration of linguistic clues, allowing us to understand how effectively the evaluated model can expand traditional solution paths and uncover new insights within a complex search space; (3) Creative Writing (Yao et al. 2023), a task to construct a coherent passage with four paragraphs, each ending with one of four given sentences.This task compels LLMs to generate imaginative text with logically sound and contextually rich. Detailed information is provided in Appendix.
Overall Performance
The overall performance is reported in Table 1 and Figure 3. We can see our TSE consistently outperforms compared baseline methods across different tasks on most of the metrics. The effectiveness of TSE is attributed to expanding thought structures and exploring solution space to contribute to different aspects of reasoning. Then, we analyze each task separately as follows.
Game of 24. As shown in Table 1, the TSE-refined CoT-SC method significantly outperforms other thought structures, including vanilla CoT-SC, with an improvement of $58.56\%$, ToT by $40.50\%$, and RATT by $79.04\%$. Additionally, our enhancement of the basic CoT-SC structure achieves accuracy on GPT-4o-mini that matches the performance of the more complex ToT ($b=5$) implemented on GPT-4 (Yao et al. 2024). These results highlight the substantial improvements in reasoning capabilities brought about by our method.
Mini Crossword. In this task, we evaluate performance based on two metrics: the proportion of correct letters (out of 25 per game) and the proportion of successfully solved games. As shown in Figure 3(b), the CoT-SC with TSE achieves an impressive accuracy rate of $82.4\%$, significantly outperforming vanilla CoT-SC (by $30.2\%$), surpassing ToT by $6.9\%$. However, for the Game metric, RATT performs slightly better than our method, as it can leverage external knowledge through RAG. Specifically, vanilla CoT-SC suffers from limited exploration of thought paths, which reduces its ability to consistently generate accurate answers. These results show that TSE enhances the model’s depth and diversity in exploring potential solutions and generates more accurate and coherent outputs.
Creative Writing. As shown in Table 1, the evaluation criteria included Soundness, Innovation, Content Coherence of Reasoning, and Clarity of Expression, applying consistently across all evaluations. We rate each metric on a scale from 0 to 10 in 0.5-point increments, and calculate the overall performance score as the average of these four dimensions. A detailed figure can be found in the Appendix.
Ablation Study
Key Node Selection.
To validate the importance of key node selection, we compare (1) the original CoT-SC, (2) the random selection of one node per chain for generation, and (3) our key node selection method. As shown in Table 3, our method significantly improves on Game of 24 compared to original and random approaches. In Creative Writing, particularly for innovation, our method also demonstrates notable enhancements, proving its effectiveness in selecting nodes for generation. The superior performance of our key node selection method is primarily due to its ability to identify nodes containing crucial information for problem-solving. The model only utilizes the most influential information and thus avoids redundancy, thus enhancing the model’s efficiency and creativity.
Connection Node Selection.
We then test the necessity of connection node selection by comparing (1) randomly selecting key nodes to connect, (2) selecting the key nodes based on lower inference layers, and (3) our method. As shown in Table 3, our method is significantly superior to random selection and layer-based selection on overall reasoning performance, particularly in coherence. This is because random selection lacks specificity and cannot guarantee that the chosen nodes will effectively support the content of new nodes. Although layer-based selection considers the structural hierarchy of information, it does not necessarily reflect the actual importance or applicability of the information. Our method analyzes relationships and reasoning paths between nodes, thus maintaining and strengthening the information flow’s coherence and depth.
Collaborative Reasoning.
In this experiment, we test the importance of the collaboration method in TSE. We compare four collaboration methods: (1) aggregating answers through majority voting, (2) randomly sampling partial thought chains, (3) using only the outcomes from newly generated chains, and (4) implementing the TSE collaborative method. As shown in Table 3, TSE consistently outperforms other methods in both tasks. Meanwhile, using only new chains also achieves a high success rate in Game of 24, demonstrating their potential to offer innovative solutions. However, since these new chains lack integration with original chain information, their performance is still weaker than that of TSE’s comprehensive collaboration methods. In Creative Writing, using only new chains proved better than majority voting. This suggests that while new chains can offer innovative content, they may lack coherence and depth without sufficient integration with the original chains.
Method GPT Llama
Success Rate (%) Overall Coherence Success Rate (%) Overall Coherence
Key Node Selection
Original 13.3 5.12 5.03 5.33 4.68 4.85
Random 51.5 5.20 4.97 3.95 4.54 4.51
TSE 74.0 6.03 5.74 18 75 5.09 5.05
Connection Node Selection
Random 62.8 5.91 5.66 9.52 4.82 4.85
Layer-based 67.3 5.95 5.69 6.12 4.84 4.88
TSE 74.0 6.03 5.74 18.75 5.09 5.05
Collaborative Reasoning
Majority-Vote 46.7 5.81 5.87 10.0 4.85 4.85
Random Sampling 27.3 5.90 5.89 12.5 4.92 4.90
New Chains Only 64.0 5.95 5.87 18.06 5.11 5.06
TSE 74.0 6.03 5.90 18.75 5.09 5.05
Table 2: Impact of node selection methods.
Strategy Analysis
Impact of Key Node Selection Methods.
In this experiment, we compare the impact of two key node selection methods—gradient-based and self-prompting—on Llama’s CoT-SC reasoning performance. As shown in Figure 4, both methods achieve comparable results in the Game of 24, while self-prompting performs better in Creative Writing. This indicates both approaches effectively select critical nodes, but self-prompting additionally helps the model revisit overlooked thought connections and blind spots. Consequently, the model generates more reliable information and produces more creative content. Thus, for tasks demanding higher creativity, self-prompting can encourage the model to move beyond conventional thought patterns and foster innovative solutions.
Game of 24 Creative Writing
Method Success Rate (%) Overall Coherence
Connection Node Selection Methods
Gradient 13.61 4.88 4.86
Relevance 16.32 4.91 4.94
Collaborative Reasoning Methods
CWS 12.93 4.97 5.01
LM-as-a-Judge 18.05 5.09 5.05
Table 3: Comparison of connection node selection and collaborative reasoning strategy.
(a) Performance by letter.
(b) Performance by game.
Figure 3: Performance of GPT-4o-mini on Mini CrossWord.
Figure 4: Performance of key node selection methods.
Comparison of Connection Node Selection Methods.
We then compare gradient comparison and semantic relevance for selecting connection nodes when generating new reasoning paths on Llama. As shown in Table 3, semantic relevance selection achieves better performance on both Game of 24 and coherence in Creative Writing. The difference possibly arises from two main reasons: (1) Nodes with larger relative gradients do not necessarily have a stronger influence on the content when generating new nodes. (2) Compared to gradient-based selection, the semantic relevance method more effectively captures the actual content and meaning relationships between nodes, leading to more suitable connection points for new nodes. This demonstrates that considering semantic relevance when selecting connection points provides more effective support for reasoning performance.
Comparison of Collaborative Reasoning Methods.
In this experiment, we compare the performance of CWS and LM-as-a-Judge. As shown in Table 3, in task Game of 24, which requires dense logic and has well-defined objectives, the LM-as-a-Judge method accurately and quickly identifies the correct answers by directly evaluating each thought chain and selecting the optimal solution. This method relies on the model’s ability to assess individual outputs, allowing for rapid selection of the best solution—especially useful in tasks that require precise calculations and quick responses. In task Creative Writing, which requires innovation and diversity, CWS evaluates and weights the contributions of each thought chain. The final output represents a blend of multiple perspectives and enhances the richness and depth of the narrative significantly. We further analyze the efficacy and provide a case study to illustrate the effectiveness of TSE generation and collaboration in the Appendix.
Conclusion
In this study, we introduce TSE, a novel approach to enhance the reasoning structures of LLMs. TSE generates new thought branches based on existing thought paths to explore previously overlooked solutions. The generated new reasoning nodes and chains are incorporated into thought structures to explore diverse reasoning directions in terms of a reasoning task. Our experiments across multiple reasoning datasets demonstrate the effectiveness of the TSE. The detailed analysis reveals the utilization of each component in TSE during the integration of diverse thought processes.
Appendix A Experimental Settings
Baselines
We apply our method to the simplest multi-chain thought structure and compare the results with several baseline methods, including CoT, Vanilla CoT-SC, ToT, and RATT. This comparison aims to illustrate how our method enhances the thought structure compared to existing approaches. For a consistent evaluation, we standardize the depth to five across all methods. Specifically, we set the number of chains to $k=5$ and generate $b=5$ candidates at each step for CoT-SC and ToT, respectively. For RATT, the model generates and integrates five candidate results at each decision point and uses Wikipedia [^1] https://en.wikipedia.org as the external resource.
Task Description
We evaluate TSE and the baseline methods on three reasoning datasets with specific tasks: (1) Game of 24, a mathematical challenge whose objective is to use the four basic arithmetic operations to make four given numbers equal 24. The task requires the language models to combine multiple operations to achieve a target outcome, which can evaluate arithmetic reasoning and logical problem-solving capabilities. (2) Mini Crosswords, a game of 5×5 mini crosswords and each input includes the 5 horizontal and 5 vertical clues. The expected output is a completed $5×5$ crossword board containing 25 letters. To solve this task, the model requires deeper exploration and strategic integration of linguistic clues, allowing us to understand how effectively the evaluated model can expand traditional solution paths and uncover new insights within a complex search space; (3) Creative Writing, a task to construct a coherent passage with four paragraphs, each ending with one of four given sentences.This task compels LLMs to generate imaginative text withlogically sound, and contextually rich. To ensure a rigorous and comprehensive assessment, we conduct half of the evaluations using GPT-4 (OpenAI 2023), while the other half are performed by a panel of expert human annotators. In this task, we select 100 open-ended questions from several prompts listed on Reedsy. com [^2] https://blog.reedsy.com/creative-writing-prompts/“#:“˜:text=When“%20the“%20idea“%20to“%20start“%20a“%20weekly“%20newsletter as our input.
Thought Chain Efficacy
In this experiment, we conduct a detailed comparison between the success rates of original thought chains and new chains generated by our method in the Game of 24 task. Experimental results indicate that for each problem, the success rate of having at least one correct answer in the original CoT-SC is approximately $50.7\%$, while the new chains generated by our method achieve a success rate of $64.0\%$. Moreover, despite improving accuracy, the overlap of problems successfully solved by both the old and new chains is only $35.3\%$. This low level of overlap suggests that our method effectively explores new areas that the original thought structures do not address and brings a significant increase in the overall success rate.
Creative Writing
In Figure 5 we provide detailed results on task Creative Writing. TSE method outperforms other methods in three aspects and is only slightly weaker than RATT with RAG.
Figure 5: Performance of different methods on GPT-4o-mini for task Creative Writing.
Appendix B Detailed Information of Case Study
Case Analysis
As our approach significantly enhances the reasoning capabilities of thought structures while maintaining minimal computational cost, it effectively addresses gaps within existing structures. Figure 6 and Table 4 provides a specific example of this enhancement by comparing the original thought chains with the new chains generated by our method. In the provided case study, we see the reasoning prompt “Explain why it is important for children to learn mathematics”. The original thought chains (Thought Chain 1 and Thought Chain 2) are linear and somewhat limited in scope, focusing and repeating on foundational aspects like “Critical Thinking”, “Academic Success” and “Career Opportunities”. However, the new chain generated by our method explores additional dimensions, such as ‘‘Enhancement of Cognitive and Memory Skills” and “Preparation for the Digital Age”. These new perspectives enrich the argument and provide a more comprehensive understanding of the importance of mathematics.
Figure 6: Case study of comparison of original and new chains.
Thought Chain 1 Thought Chain 2 New Chain
Foundation for Critical Thinking: Mathematics fosters critical and logical thinking, enhancing problem-solving and analytical skills necessary for daily life. Foundational Skills: Mathematics equips children with essential daily skills like counting, measuring, and problem-solving, crucial for tasks such as budgeting, shopping, and time management. Career Opportunities: Proficiency in mathematics is crucial for careers in engineering, finance, data analysis, and technology, preparing children for diverse job prospects in a competitive market.
Real-World Applications: Mathematics is crucial for managing daily activities like finances, cooking, and scheduling, empowering children with the confidence to handle them effectively. Cognitive Development: Math enhances logical reasoning, critical thinking, and problem-solving abilities, benefiting other subjects and real-life situations. Academic Success: A strong foundation in mathematics is essential for excelling across subjects like science, technology, engineering, and social studies, enabling access to advanced learning opportunities.
Academic Success: A strong mathematics foundation is essential for excelling in subjects like science, technology, engineering, and social studies, paving the way for advanced educational opportunities. Academic Success: Proficiency in mathematics underpins success in STEM fields, crucial for today’s job market, and opens doors to advanced educational opportunities. Foundational Skills for Future Success: Mastering mathematics early equips children with critical thinking and problem-solving skills essential for academic and professional success in fields like science, technology, and engineering, preparing them for a competitive global environment.
Career Opportunities: Proficiency in mathematics is key for careers in engineering, finance, technology, medicine, and data science, equipping children for diverse job prospects in a competitive market. Career Opportunities: Proficiency in mathematics is crucial for careers in engineering, finance, data analysis, and technology, preparing children for a broad range of opportunities in a competitive job market. Enhancement of Cognitive and Memory Skills: Regular math practice boosts cognitive skills like attention to detail, memory retention, and mental arithmetic, essential for efficient learning and problem-solving in academic and everyday life.
Development of Persistence and Resilience: Mathematics teaches children systematic problem-solving, a skill invaluable in academics and beyond. Problem-Solving Skills: Mathematics teaches children systematic problem-solving, a skill invaluable in academics and beyond. Preparation for the Digital Age: Mathematics prepares children for the digital world by teaching them to understand and manipulate numerical data, essential skills for coding and navigating technological advancements.
Table 4: Reasoning Prompt: Explain why it is important for children to learn mathematics.